OpenStudy (anonymous):

suppose (-6,2) is a point on the graph of y=g(x). what point is on the graph of y=-4g(x-7)+1

OpenStudy (anonymous):

if x = 1 you get \[-4\times g(1-7)+1=-4\times g(-6)+1=-4\times 2+1=-8+1=-7\]

myininaya (myininaya):

if g(-6)=2 then if x-7=-6 >x=1 so y=-4g(1-7)+1=-4g(-6)+1=-4(2)+1=-8+1=-7

OpenStudy (anonymous):

so \[(1,-7)\] is on the graph of \[-4g(x-7)+1\]

OpenStudy (anonymous):

thank you